OpenStudy (anonymous):

b+2/9 = 1 3/5 PLEASE HELPPPP WHATS B?

OpenStudy (anonymous):

@Yolo? HELP ME!!!!!!!!!!!!

OpenStudy (anonymous):

YOUR SMART

OpenStudy (compassionate):

\[b + \frac{ 2 }{ 9 } = 1\frac{ 3 }{ 5}\] Is that right?

OpenStudy (anonymous):

Thank you so muchooooooo!

OpenStudy (compassionate):

b = 62/5 :P

OpenStudy (compassionate):

I was asking if I read the equation right.

OpenStudy (anonymous):

yeah thats right

OpenStudy (compassionate):

so b = 62/5

OpenStudy (compassionate):

Wait. B = 62/45 ^^^ That is correct!

OpenStudy (anonymous):

wait its not on my answer choice

OpenStudy (compassionate):

What are your choices?

OpenStudy (anonymous):

look it says this... Solve 27 viviansalem b+2/9 = 1 3/5

OpenStudy (compassionate):

Can you write the equation using the draw tool at the bottom? I'm having trouble understanding the format it's in.  ^ Is that correct?